# Exposing Vulnerabilities in Explanation for Time Series Classifiers via Dual-Target Attacks

### v1

@@ 1 Introduction
Deep neural networks (DNNs) have achieved success in time series classification, enabling breakthroughs in healthcare, finance, and industrial systems (Wu et al., 2021; Zhou et al., 2022; Liu et al., 2023b; Wang et al., 2024; Liu et al., 2025; Wen et al., 2024; Zhou et al., 2023; Mohammadi Foumani et al., 2024). Yet, their black-box nature leaves a fundamental gap: While we can train models that classify signals with high accuracy, we often cannot tell why they made a specific decision. This opacity is especially concerning in high-stakes scenarios such as medical diagnosis or fault detection, where practitioners must understand the evidence behind a prediction before acting on it (Rojat et al., 2021; Theissler et al., 2022). To bridge this gap, interpretable time series deep learning systems (ITSDLS) pair a predictor with an explainer that highlights salient time–channel regions deemed responsible for the prediction (Queen et al., 2023; Liu et al., 2024). In practice, such explanations are increasingly used for model debugging, expert verification, and even adversarial monitoring, providing an appealing layer of human oversight (Rojat et al., 2021; Šimić et al., 2021).
Figure 1: Targeted manipulation of prediction and explanation on an ECG sample. Top (Target Setup): The input is correctly classified as Normal ($\hat{y}=0$). The green band defines the reference explanation $\mathbf{A}^{\prime}$ that the attacker aims to fabricate. Bottom (TSEF Attack): The proposed TSEF attack successfully flips the model’s prediction to Abnormal ($\hat{y}=1$). Crucially, the resulting explanation (generated by TimeX++ (Liu et al., 2024)) concentrates on the target region $\mathbf{A}^{\prime}$ defined above.
@@ 2 Related Work
Time Series Explainers. While Explainable AI (XAI) has been extensively studied for image and text (Danilevsky et al., 2020; Samek et al., 2021), explaining time series models presents unique challenges due to complex temporal dependencies (Rojat et al., 2021; Theissler et al., 2022). Early works adapted generic attribution methods (e.g., Integrated Gradients) to the temporal domain (Sundararajan et al., 2017; Lundberg and Lee, 2017; Ismail et al., 2020). Recent approaches develop time series specific explainers, such as perturbation-based methods (Crabbé and Van Der Schaar, 2021; Leung et al., ; Chuang et al., ; Huang et al., 2025) and mask-based methods like TimeX and TimeX++ (Queen et al., 2023; Liu et al., 2024), which produce in-distribution and coherent explanations. Our focus is orthogonal: we study the security of these explainers under adaptive adversaries, showing that an attacker can enforce a target label while steering the explainer toward a reference explanation.
@@ 3 Preliminaries and Problem Definition
Time Series Explanations. An explanation method, or explainer, $\mathcal{H}^{E}$, is a function that computes the importance of each feature at each time step for a prediction. For an input $\mathbf{X}$, the explainer outputs a saliency map $\mathbf{A}\in\mathbb{R}^{T\times D}$, where the magnitude of each element $\mathbf{A}[t,d]$ corresponds to the importance of the feature $d$ at time $t$. This general definition encompasses a wide range of explainers (Sundararajan et al., 2017; Crabbé and Van Der Schaar, 2021; Huang et al., 2025; Leung et al., ; Chuang et al., ), including TimeX (Queen et al., 2023) and TimeX++ (Liu et al., 2024).
@@ 5.1 Experimental Setup
Datasets and Evaluation metrics. We evaluate on six benchmarks: three synthetic datasets (LowVar, SeqComb-UV, SeqComb-MV) and three real-world datasets (ECG (Moody and Mark, 2001), PAM (Reiss and Stricker, 2012), Epilepsy (Andrzejak et al., 2001)). We use the train–test splits from prior work (Queen et al., 2023; Liu et al., 2024) and attack correctly classified test samples. Dataset details are in Appendix E. For explanations, we report AUPRC, AUR, and AUP to quantify alignment between the produced importance maps and reference explanations. For classification, we report F1 score and targeted Attack Success Rate (ASR), where ASR is the fraction of attacked samples whose predicted label equals the randomly chosen target label $y^{\prime}$.
@@
Threat models: (1) Time series interpreters. We adopt TimeX++ (Liu et al., 2024), TimeX (Queen et al., 2023), and Integrated Gradients (IG) (Sundararajan et al., 2017) as the time series explanation methods under attack. (2) Time series classifier. We use a transformer-based model (Vaswani et al., 2017) as the time series classifier for all datasets. Baseline attacks and settings. We compare against local/global Gaussian perturbations (Nguyen et al., 2024), Random Sign Attack, BlackTreeS (Ding et al., 2023), SFAttack (Gu et al., 2025), PGD (Madry et al., 2017), and ADV2 (Zhang et al., 2020), a representative joint attack originally designed for image classifiers and explainers. We adapt baselines to the targeted setting by optimizing a targeted classification toward $y^{\prime}$; To adapt ADV2 to time series, we replace its interpretation loss with a time series explanation-alignment term toward $\mathbf{A}^{\prime}$, so that all methods are evaluated under the same targeted objectives as TSEF. All attacks use an $\ell_{\infty}$ budget $\epsilon=0.1$. To handle heterogeneous signal scales, we use an instance-wise bound $\epsilon^{\prime}=\epsilon\cdot(\max(\mathbf{X})-\min(\mathbf{X}))$ and enforce $\|\mathbf{X}-\tilde{\mathbf{X}}\|_{\infty}\leq\epsilon^{\prime}$ via projection at each step (Alg. 1). Hyperparameters are tuned on validation sets; full settings are in Appendix F.
@@ Appendix A Related Work Details
Time series explainers.
The increasing deployment of deep models in high-stakes applications has made human-understandable explanations a key requirement for trustworthy decision making. While the broader XAI literature is dominated by vision and language settings (Danilevsky et al., 2020; Madsen et al., 2022; Bodria et al., 2023; Linardatos et al., 2020; Samek et al., 2021; Buhrmester et al., 2021), time series interpretability has received comparatively less attention despite its practical importance (Rojat et al., 2021; Theissler et al., 2022). Early efforts often adapt generic attribution tools such as Integrated Gradients and SHAP to temporal inputs (Sundararajan et al., 2017; Lundberg and Lee, 2017). However, time series poses modality-specific challenges: discriminative cues may be embedded in long-term trends, feature relevance can vary over time and across channels, and common perturbation-based explainers may produce out-of-distribution counterfactuals that compromise faithfulness (Liu et al., 2024; Ye and Keogh, 2009; Szegedy et al., 2013). To systematically assess these issues, Ismail et al. (2020) benchmark saliency methods for multivariate time series on synthetic data with ground-truth temporal–feature importance, finding that many explainers are unreliable across RNN/TCN/Transformer backbones; they also propose Temporal Saliency Rescaling (TSR) to improve localization by first identifying salient time steps and then refining per-feature attributions within them. More recent explainers explicitly target in-distribution and pattern-coherent rationales. TimeX (Queen et al., 2023) encourages consistency between explainer and model latent representations, and TimeX++ (Liu et al., 2024) further generates label-preserving temporal patterns via an information-theoretic objective that minimizes mutual information with the original series while maximizing mutual information with the prediction. Our work is complementary but security-focused: rather than proposing another explainer, we study whether such modern, pattern-aware explainers remain reliable under adaptive adversaries. We show that an attacker can simultaneously force a target label and steer explanations toward a chosen reference rationale, enabling a realistic cover-up failure mode for explanation-based auditing.
Adversarial attacks on time series Models
@@
Explainer robustness in vision and NLP.
Compared to time series, explainer robustness has been more extensively studied in vision and NLP. In image classification, Ghorbani et al. (2019) show that saliency maps from popular gradient-based methods can be drastically altered by imperceptible perturbations while preserving the predicted label, revealing the fragility of attribution-based explanations. Zhang et al. (2020) further demonstrate black-box attacks on interpretable systems, including attacks that selectively fool explanations or jointly mislead predictions and explanation reports. In NLP, Ivankay et al. (2022) propose TextExplanationFooler (TEF), using semantics-preserving textual perturbations to significantly change token-level attributions while keeping classifier outputs stable. While these results establish that explanation manipulation is possible in other modalities, transferring the same capability to time series is not straightforward. time series auditing often relies on temporally coherent, pattern-like rationales, and our analysis highlights a modality-specific high-dimensional paradox: although the $T\times D$ input space offers a large surface for flipping predictions, unconstrained optimization tends to yield dense, temporally incoherent updates that are difficult to steer toward a target explanation (Section 4.1). Our contribution is time series specific: we instantiate a time–frequency structured dual-target attack that achieves pattern-level control, enabling targeted label manipulation while matching a reference rationale for explanation-based auditing.
@@
For standard time series explainers, the attribution $A_{i}(\mathbf{X})$ is a complex, non-linear function of the input (Sundararajan et al., 2017; Crabbé and Van Der Schaar, 2021; Huang et al., 2025; Leung et al., ; Chuang et al., ; Queen et al., 2023; Liu et al., 2024). Unless an explainer explicitly employs a hard-masking mechanism on $\Omega^{c}$, the gradient $\nabla A_{i}(\mathbf{X})$ is non-vanishing for the majority of off-window coordinates (Ghorbani et al., 2019; Dombrowski et al., 2019). Furthermore, as the classification gradient sign $\mathrm{sign}(g_{c})\in\{\pm 1\}^{d}$ is a dense vector, the directional derivative $\nabla A_{i}(\mathbf{X})^{\top}\mathrm{sign}(g_{c})$ would vanish only under exact, non-generic cancellation across many input dimensions. Such perfect cancellation is statistically improbable and unstable in the presence of even infinitesimal noise, ensuring that $\gamma$ remains bounded away from zero in practice.
@@ Appendix E Description of Datasets
We follow the dataset construction and preprocessing protocols of (Queen et al., 2023; Liu et al., 2024) and use both synthetic and real-world benchmarks. Below we summarize the datasets used in our experiments and how ground-truth explanations are defined when available.
@@ E.1 Synthetic datasets
We adopt three synthetic datasets with known ground-truth explanations to study how well attacks and explainers can identify the underlying predictive signal. Following (Queen et al., 2023; Liu et al., 2024), all synthetic time series are built on top of a non-autoregressive moving average (NARMA) noise process, into which class-specific patterns are inserted. The design follows standard recommendations for synthetic benchmarks that avoid shortcut solutions and require the model to detect precise temporal patterns rather than simple heuristics. For each synthetic dataset, we generate 5,000 training samples, 1,000 test samples, and 100 validation samples.

### v2

@@ 1 Introduction
Deep neural networks (DNNs) have achieved success in time series classification, enabling breakthroughs in healthcare, finance, and industrial systems (Liu et al., 2025b; Wen et al., 2024; Zhou et al., 2023; Mohammadi Foumani et al., 2024). Yet, their black-box nature leaves a fundamental gap: While we can train models that classify signals with high accuracy, we often cannot tell why they made a specific decision. This opacity is especially concerning in high-stakes scenarios such as medical diagnosis or fault detection, where practitioners must understand the evidence behind a prediction before acting on it (Rojat et al., 2021; Theissler et al., 2022). To bridge this gap, interpretable time series deep learning systems (ITSDLS) pair a predictor with an explainer that highlights salient time–channel regions deemed responsible for the prediction (Queen et al., 2023; Liu et al., 2024b). In practice, such explanations are increasingly used for model debugging, expert verification, and even adversarial monitoring, providing an appealing layer of human oversight (Rojat et al., 2021; Šimić et al., 2021).
Figure 1: Targeted manipulation of prediction and explanation on an ECG sample. Top (Target Setup): The input is correctly classified as Normal ($\hat{y}=0$). The green band defines the reference explanation $\mathbf{A}^{\prime}$ that the attacker aims to fabricate. Bottom (TSEF Attack): The proposed TSEF attack successfully flips the model’s prediction to Abnormal ($\hat{y}=1$). Crucially, the resulting explanation (generated by TimeX++ (Liu et al., 2024b)) concentrates on the target region $\mathbf{A}^{\prime}$ defined above.
@@ 2 Related Work
Time Series Explainers. While Explainable AI (XAI) has been extensively studied for image and text (Danilevsky et al., 2020; Samek et al., 2021), explaining time series models presents unique challenges due to complex temporal dependencies (Rojat et al., 2021; Theissler et al., 2022). Early works adapted generic attribution methods (e.g., Integrated Gradients) to the temporal domain (Sundararajan et al., 2017; Lundberg and Lee, 2017; Ismail et al., 2020). Recent approaches develop time series specific explainers, such as perturbation-based methods (Crabbé and Van Der Schaar, 2021; Leung et al., ; Chuang et al., ; Huang et al., 2025) and mask-based methods like TimeX and TimeX++ (Queen et al., 2023; Liu et al., 2024b), which produce in-distribution and coherent explanations. Our focus is orthogonal: we study the security of these explainers under adaptive adversaries, showing that an attacker can enforce a target label while steering the explainer toward a reference explanation.
@@ 3 Preliminaries and Problem Definition
Time Series Explanations. An explanation method, or explainer, $\mathcal{H}^{E}$, is a function that computes the importance of each feature at each time step for a prediction. For an input $\mathbf{X}$, the explainer outputs a saliency map $\mathbf{A}\in\mathbb{R}^{T\times D}$, where the magnitude of each element $\mathbf{A}[t,d]$ corresponds to the importance of the feature $d$ at time $t$. This general definition encompasses a wide range of explainers (Sundararajan et al., 2017; Crabbé and Van Der Schaar, 2021; Huang et al., 2025; Leung et al., ; Chuang et al., ), including TimeX (Queen et al., 2023) and TimeX++ (Liu et al., 2024b).
@@ 5.1 Experimental Setup
Datasets and Evaluation metrics. We evaluate on six benchmarks: three synthetic datasets (LowVar, SeqComb-UV, SeqComb-MV) and three real-world datasets (ECG (Moody and Mark, 2001), PAM (Reiss and Stricker, 2012), Epilepsy (Andrzejak et al., 2001)). We use the train–test splits from prior work (Queen et al., 2023; Liu et al., 2024b) and attack correctly classified test samples. Dataset details are in Appendix E. For explanations, we report AUPRC, AUR, and AUP to quantify alignment between the produced importance maps and reference explanations. For classification, we report F1 score and targeted Attack Success Rate (ASR), where ASR is the fraction of attacked samples whose predicted label equals the randomly chosen target label $y^{\prime}$.
@@
Threat models: (1) Time series interpreters. We adopt TimeX++ (Liu et al., 2024b), TimeX (Queen et al., 2023), and Integrated Gradients (IG) (Sundararajan et al., 2017) as the time series explanation methods under attack. (2) Time series classifier. We use a transformer-based model (Vaswani et al., 2017) as the time series classifier for all datasets. Baseline attacks and settings. We compare against local/global Gaussian perturbations (Nguyen et al., 2024), Random Sign Attack, BlackTreeS (Ding et al., 2023), SFAttack (Gu et al., 2025), PGD (Madry et al., 2017), and ADV2 (Zhang et al., 2020), a representative joint attack originally designed for image classifiers and explainers. We adapt baselines to the targeted setting by optimizing a targeted classification toward $y^{\prime}$; To adapt ADV2 to time series, we replace its interpretation loss with a time series explanation-alignment term toward $\mathbf{A}^{\prime}$, so that all methods are evaluated under the same targeted objectives as TSEF. All attacks use an $\ell_{\infty}$ budget $\epsilon=0.1$. To handle heterogeneous signal scales, we use an instance-wise bound $\epsilon^{\prime}=\epsilon\cdot(\max(\mathbf{X})-\min(\mathbf{X}))$ and enforce $\|\mathbf{X}-\tilde{\mathbf{X}}\|_{\infty}\leq\epsilon^{\prime}$ via projection at each step (Alg. 1). Hyperparameters are tuned on validation sets; full settings are in Appendix F.
@@ Appendix A Related Work Details
Time series analysis and time series explainers.
Time series analysis plays a foundational role across a wide range of domains such as finance (Sezer et al., 2020; Emami Gohari et al., 2024), traffic systems (Wu et al., 2021; Zhou et al., 2022; Liu et al., 2023b, 2025a), healthcare (Lauritsen et al., 2020; Talukder et al., 2025), epidemiology (Liu et al., 2024a; Wan et al., 2025), audio processing (Baevski et al., 2020; Wang et al., 2023) etc. The increasing deployment of deep models in high-stakes applications has made human-understandable explanations a key requirement for trustworthy decision making. While the broader XAI literature is dominated by vision and language settings (Danilevsky et al., 2020; Madsen et al., 2022; Bodria et al., 2023; Linardatos et al., 2020; Samek et al., 2021; Buhrmester et al., 2021), time series interpretability has received comparatively less attention despite its practical importance (Rojat et al., 2021; Theissler et al., 2022). Early efforts often adapt generic attribution tools such as Integrated Gradients and SHAP to temporal inputs (Sundararajan et al., 2017; Lundberg and Lee, 2017). However, time series poses modality-specific challenges: discriminative cues may be embedded in long-term trends, feature relevance can vary over time and across channels, and common perturbation-based explainers may produce out-of-distribution counterfactuals that compromise faithfulness (Liu et al., 2024b; Ye and Keogh, 2009; Szegedy et al., 2013). To systematically assess these issues, Ismail et al. (2020) benchmark saliency methods for multivariate time series on synthetic data with ground-truth temporal–feature importance, finding that many explainers are unreliable across RNN/TCN/Transformer backbones; they also propose Temporal Saliency Rescaling (TSR) to improve localization by first identifying salient time steps and then refining per-feature attributions within them. More recent explainers explicitly target in-distribution and pattern-coherent rationales. TimeX (Queen et al., 2023) encourages consistency between explainer and model latent representations, and TimeX++ (Liu et al., 2024b) further generates label-preserving temporal patterns via an information-theoretic objective that minimizes mutual information with the original series while maximizing mutual information with the prediction. Our work is complementary but security-focused: rather than proposing another explainer, we study whether such modern, pattern-aware explainers remain reliable under adaptive adversaries. We show that an attacker can simultaneously force a target label and steer explanations toward a chosen reference rationale, enabling a realistic cover-up failure mode for explanation-based auditing.
Adversarial attacks on time series models.
@@
Explainer robustness in other domains.
Compared to time series, explainer robustness has been more extensively studied in vision and NLP. In image classification, Ghorbani et al. (2019) show that saliency maps from popular gradient-based methods can be drastically altered by imperceptible perturbations while preserving the predicted label, revealing the fragility of attribution-based explanations. Zhang et al. (2020) further demonstrate black-box attacks on interpretable systems, including attacks that selectively fool explanations or jointly mislead predictions and explanation reports. In NLP, Ivankay et al. (2022) propose TextExplanationFooler (TEF), using semantics-preserving textual perturbations to significantly change token-level attributions while keeping classifier outputs stable. Beyond these, attacks on explainers have also been explored in the graph domain (Fan et al., 2023; Li et al., 2024). While these results establish that explanation manipulation is possible in other modalities, transferring the same capability to time series is not straightforward. time series auditing often relies on temporally coherent, pattern-like rationales, and our analysis highlights a modality-specific high-dimensional paradox: although the $T\times D$ input space offers a large surface for flipping predictions, unconstrained optimization tends to yield dense, temporally incoherent updates that are difficult to steer toward a target explanation (Section 4.1). Our contribution is time series specific: we instantiate a time–frequency structured dual-target attack that achieves pattern-level control, enabling targeted label manipulation while matching a reference rationale for explanation-based auditing.
@@
For standard time series explainers, the attribution $A_{i}(\mathbf{X})$ is a complex, non-linear function of the input (Sundararajan et al., 2017; Crabbé and Van Der Schaar, 2021; Huang et al., 2025; Leung et al., ; Chuang et al., ; Queen et al., 2023; Liu et al., 2024b). Unless an explainer explicitly employs a hard-masking mechanism on $\Omega^{c}$, the gradient $\nabla A_{i}(\mathbf{X})$ is non-vanishing for the majority of off-window coordinates (Ghorbani et al., 2019; Dombrowski et al., 2019). Furthermore, as the classification gradient sign $\mathrm{sign}(g_{c})\in\{\pm 1\}^{d}$ is a dense vector, the directional derivative $\nabla A_{i}(\mathbf{X})^{\top}\mathrm{sign}(g_{c})$ would vanish only under exact, non-generic cancellation across many input dimensions. Such perfect cancellation is statistically improbable and unstable in the presence of even infinitesimal noise, ensuring that $\gamma$ remains bounded away from zero in practice.
@@ Appendix E Description of Datasets
We follow the dataset construction and preprocessing protocols of (Queen et al., 2023; Liu et al., 2024b) and use both synthetic and real-world benchmarks. Below we summarize the datasets used in our experiments and how ground-truth explanations are defined when available.
@@ E.1 Synthetic datasets
We adopt three synthetic datasets with known ground-truth explanations to study how well attacks and explainers can identify the underlying predictive signal. Following (Queen et al., 2023; Liu et al., 2024b), all synthetic time series are built on top of a non-autoregressive moving average (NARMA) noise process, into which class-specific patterns are inserted. The design follows standard recommendations for synthetic benchmarks that avoid shortcut solutions and require the model to detect precise temporal patterns rather than simple heuristics. For each synthetic dataset, we generate 5,000 training samples, 1,000 test samples, and 100 validation samples.